OpenStudy (anonymous):

Four employees are cleaning out four different wishing wells at an amusement park. Tracy collects $5.80 more than Asuk. Kyle collects half as much as racy. Beverly collects 3 times as much as Kyle. How much money does each employee collect if the total amount of money collected is $246.20

OpenStudy (tkhunny):

Please NAME the amounts. Then the problem statement can be translated.

OpenStudy (anonymous):

we have to introduce a variable we could use say \(x\)

OpenStudy (anonymous):

and we also have to say what it represents. i will say let \(x\) be the amount of money Asuk (?!) collects

OpenStudy (anonymous):

that means, that since tracy collects $5.80 more, if Asuk collects \(x\) dollars, how much does Tracy collect?

OpenStudy (anonymous):

if it is not clear, let me know and i will tell you

OpenStudy (anonymous):

x+5.80

OpenStudy (anonymous):

ok good so now we can go on, because the method is all the same

OpenStudy (anonymous):

Asuk collects \(x\) Tracy collects \(x+5.8\) Kyle collects ?

OpenStudy (anonymous):

kyle is the confusing one but I believe it's 1/2 (x+5.80)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or \(\frac{x+5.8}{2}\) but the way you wrote it is probably better

OpenStudy (anonymous):

ok now do I add all the numbes and the equal 246.20?

OpenStudy (anonymous):

sorry numbers

OpenStudy (anonymous):

you have one more to go

OpenStudy (anonymous):

Asuk \(x\) Tracy \(x+5.8\) Kyle \(\frac{1}{2}(x+5.8)\) Beverly

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

b=3x+1/2(x+5.80)

OpenStudy (anonymous):

i think that is not quite right "3 times as much as kyle" for beverly kyle is \(\frac{1}{2}(x+5.8)\) multiply it by 3

OpenStudy (anonymous):

(3)+1/2(=5.8)

OpenStudy (anonymous):

times, not plus

OpenStudy (anonymous):

sorry plus 5.8

OpenStudy (anonymous):

I'm not sure how to write it sorry can you show me

OpenStudy (anonymous):

sure if i have $100 and you have three time as much, how much do you have?

OpenStudy (anonymous):

ok i don't mean to be annoying if i have $100 and you have 3 times as much, you have $300 if kyle has \(\frac{1}{2}(x+5.8)\) and beverly has three times as much, she has \[3\times \frac{1}{2}(x+5.8)\] or \[\frac{3}{2}(x+5.8)\]

OpenStudy (anonymous):

omg my internet died o sorry please help

OpenStudy (anonymous):

No you are helping me to do the problems myself, I understand I don't just want the answer but help thank you so much